OpenStudy (anonymous):

Find the exact value of tan 22.5.

OpenStudy (anonymous):

Use the half-angle identity tan(x/2) = ±sin(x)) / (1 + cos(x)) The angle is in the first quadrant, so its tangent is positive. sin(45°) = cos(45°) = √2 / 2 If you rationalize the denominator and simplify the result, you should get √2 - 1